OpenStudy (babyslapmafro):

Find the instantaneous rate of change of y with respect to x at an arbitrary value of x0. y=2x^2

OpenStudy (zzr0ck3r):

what is the derrivative of y = 2x^2 in respect to x?

OpenStudy (zzr0ck3r):

have you learned that yet? Or are you doing (f(x+h) - f(x))/h ?

OpenStudy (babyslapmafro):

i have not covered derivatives yet

OpenStudy (babyslapmafro):

f(x+h) - f(x))/h

OpenStudy (zzr0ck3r):

ok so have you done that with this yet?

OpenStudy (zzr0ck3r):

\[\lim_{h \rightarrow 0}\frac{ 2(x+h)^2 - 2x^2 }{ h }\]

OpenStudy (babyslapmafro):

\[\lim_{h \rightarrow0 }\frac{ 2(x _{0}+h)^{2}-2x ^{2} }{ h }\]

OpenStudy (zzr0ck3r):

weird...

OpenStudy (zzr0ck3r):

ok so simplify that and wahwt do we get?

OpenStudy (babyslapmafro):

i'm having a hard time simplifying, my algebra is pathetic

OpenStudy (zzr0ck3r):

ok first step, is to foil the (x+h)^2

OpenStudy (babyslapmafro):

2x^2+4xh+2h^2

OpenStudy (babyslapmafro):

?

OpenStudy (zzr0ck3r):

\[\frac{ 2(x^2+2xh+h^2)-2x^2 }{ h }\]

OpenStudy (babyslapmafro):

oh ok

OpenStudy (zzr0ck3r):

\[\frac{ 2x^2+4xh+2h^2 - 2x^2 }{ h }\]

OpenStudy (babyslapmafro):

ya that's where i got before, now im really stuck

OpenStudy (babyslapmafro):

although i know 2x^2 is cancelled

OpenStudy (zzr0ck3r):

simplify that top and you get what?

OpenStudy (babyslapmafro):

4xh+2h^2

OpenStudy (zzr0ck3r):

ok so we have (4xh+2h^2)/h

OpenStudy (zzr0ck3r):

this is the same as (4xh)/h + (2h^2)/h right?

OpenStudy (zzr0ck3r):

dont forget about the h on the bottom

OpenStudy (babyslapmafro):

ok, yes, i'm following you

OpenStudy (zzr0ck3r):

simplify that note: we can devide away the h, because its never 0. It just approaches 0 but never actually gets there

OpenStudy (zzr0ck3r):

so what do we have after you devide away the h in each term?

OpenStudy (babyslapmafro):

4x+4?

OpenStudy (zzr0ck3r):

not quite

OpenStudy (babyslapmafro):

4x+1?

OpenStudy (babyslapmafro):

oh nvm, just 4x

OpenStudy (zzr0ck3r):

(4xh)/h + (2h^2)/h lets look at the first term 4xh/h = ?

OpenStudy (zzr0ck3r):

correct, after you run the limit.

OpenStudy (babyslapmafro):

do you think my poor algebra skills will hold me back much in calc1?

OpenStudy (zzr0ck3r):

um, I hate to say it, but yeah. I dont think you have prro algebra skills, just do the things you know you can do. dont ever guess. You did all this by yourself pretty much.

OpenStudy (zzr0ck3r):

there is lots of algebra in calc1 and calc2

OpenStudy (zzr0ck3r):

but here in a few days or a week or so, you wont be doing this all the time. you will be able to tell me the answer to this from looking at it. what you did here way take the derivative of a function, we just went the very long way about it. there are very easy shortcuts, but you need to understand what is going on before you can use them.

OpenStudy (babyslapmafro):

I got an A in precalc and an A in trig but if i had to test-into calc1 i bet i would not be admitted into the class...i feel like something is wrong with me...

OpenStudy (babyslapmafro):

im hard-working and i never miss class and all that stuff but i dont seem to absorb anything

OpenStudy (zzr0ck3r):

nah, your fine. Everyone feels that way. You will now start to see why they tought you all that stuff. and its easier to remember once you know why you need to know it.

OpenStudy (babyslapmafro):

thanks a lot for the help...

OpenStudy (zzr0ck3r):

np, just keep doing these problems, and it will become 2nd nature I promise.

OpenStudy (babyslapmafro):

will-do